the square of a number is 6 less than 5 times the number. what is the number?
Answer by ankor@dixie-net.com(22740)

he square of a number is 6 less than 5 times the number.
what is the number?
:
n^2 = 5n -6
n^2 - 5n + 6 = 0
Factors to
(n-3)(n-2) = 0
Two solutions
n = 3
n = 2
:
You can check both solutions in the original statement